OpenStudy (anonymous):

Which point is a solution to the following system of inequalities? 2x + 3y < 6 5x – y > 10

OpenStudy (anonymous):

x=36/17 y=10/17

OpenStudy (anonymous):

The choices are a. (3,-1), b.(-1-3), c.(-1,3), d.(3,1)

OpenStudy (zepp):

A) When I played around with them. I get y > (6-2x)/3 y > 5x-10